Montserrat Volcano Observatory

Daily Report
Report for the period 16:00 26 May
to 16:00 27 May 1996

Volcanic activity at the Soufriere Hills volcano during this period has been at about the same level as yesterday. Seismicity continues to be dominated by small- to moderate-sized rockfall signals. The largest of these were at 01:59, 03:21, 07:26 and 10:56 on 27 May. These, and many of the smaller rockfalls, produced small ash plumes which drifted on the wind over the Gages and Upper Amersham areas.

There were 10 long period events and one hybrid earthquake recorded today. There has been intermittent low-amplitude broadband tremor throughout the reporting period.

There were excellent viewing conditions again today. The active areas of dome growth are still the eastern, northeastern, western and southern parts of the dome. Vigorous steaming was seen from several areas of the moat and on the dome. There were moderate amounts of blue smoke emitted from the south moat and the eastern flank of the dome. There was a new area of steam production, on the saddle between Castle Peak and the new dome; this however may have been related to rainfall last night. There was almost constant small-scale rockfall activity on the eastern and northeastern flanks of the dome. Few of the rockfalls were large, although two of them, at 9:14 and 12:07 were seen to produce small pyroclastic flows which entered the upper Tar River Valley area. There is a large highly-fractured block high on the eastern flank of the dome with a large fracture approximately 50 ft long, on its northern margin. There were almost continuous ash emissions from this fracture during a helicopter inspection in the morning. Incandescence was seen in this fracture, and in a small depression to the north of it. There was much less ash being produced from the fracture during an afternoon inspection, although the incandescence was still present.

EDM measurement were carried out on the eastern triangle to Castle Peak today. There was about 1 cm shortening in line lengths to Castle Peak since this station was last occupied on 20 May. Measurements made on the southern triangle yesterday indicate that there has been 8 to 9 mm of shortening on line lengths to Chances Peak since 21 April. The changes observed on these triangles are consistent with small scale deformation of the volcano caused by continued slow growth of the lava dome in English's Crater. The data obtained by the EDM technique is consistent with the recently concluded GPS measurements conducted by the University of Puerto Rico.

FTIR measurements were made directly from the dome again today, both from the school yard at Harris Village and from the helicopter. COSPEC measurements were made along the southern main road between Plymouth and St Patricks. The results from these measurements are currently being processed. The COSPEC data from the weekend gave estimates for the SO2 flux of about 70 and 100 tonnes per day for 25 and 26 May respectively.

During the morning helicopter inspection, estimates of the dome surface temperature were made using an optical pyrometer, a hand held thermometer which uses infra-red light to estimate temperature. The surface of the dome showed temperatures in the range 200 and 250 degrees Centigrade. The highest temperature measured was about 350 degrees.

The present level of activity at the Soufriere Hills Volcano continues to cause concern to the scientists. The MVO urges that visits to the evacuated zone are kept to a minimum. The Tar River, Long Ground and Whites areas are extremely dangerous and should not be entered under any circumstances.

Montserrat Volcano Observatory